# A Hybrid Machine Learning Framework for Predicting Hydrogen Storage Capacities: Unsupervised Feature Learning with Deep Neural Networks

### v4

@@ Abstract
In this study, we present a sophisticated hybrid machine-learning framework that significantly improves the accuracy of predicting hydrogen storage capacities in metal hydrides. This is a critical challenge due to the scarcity of experimental data and the complexity of high-dimensional feature spaces. Our approach employs the power of unsupervised learning through the use of a state-of-the-art autoencoder. This autoencoder is trained on elemental descriptors obtained from Mendeleev software, enabling the extraction of a meaningful and lower dimensional latent space from the input data. This latent representation serves as the basis for our deep multi-layer perceptron (MLP) model, which consists of five layers and shows good precision in predicting hydrogen storage capacities. Furthermore, our results show very good agreement with the results obtained with density functional theory (DFT). In addition to addressing the limitations caused by limited and unevenly distributed data in the field of hydrogen storage materials, we also focus on discovering new materials that show promising opportunities for hydrogen storage. These materials were identified using both feature-based approaches and predictions generated by a large language model. Finally, our investigation into the effectiveness of transferring weights from the autoencoder to the MLP, in addition to the latent features, suggests that while this strategy slightly improves model performance as indicated by a slightly higher R2 value and lower RMSE, it emphasizes the intricate challenge of adapting pre-trained weights for specific supervised tasks.
@@
Hydrogen storage, deep learning
@@ 1 Introduction
Climate change, energy security, and the potential depletion of resources propelled by escalating global population and technological advancement constitute pressing challenges for mankind 1, 2. To address these, renewable energy promotion is vital, with hydrogen emerging as the quintessential energy carrier for mobile and stationary applications. Hydrogen not only mitigates environmental harm but also reduces reliance on imported oil for countries devoid of natural resources 3. Various storage techniques exist for hydrogen, including pressurized gas, cryogenic liquid, and solid fuel as a chemical or physical combination with other materials like metal hydrides, complex hydrides, and carbon materials, or produced in-situ in vehicles by methanol reforming 4, 5. Each of these methods possesses unique attributes for hydrogen storage. Storage by absorption as chemical compounds or by adsorption on carbon materials carries safety benefits, necessitating a form of conversion or energy input for hydrogen release. Considerable effort has been invested into new hydrogen-storage systems, including metal, chemical, or complex hydrides, and carbon nanostructures 6.
@@
Hydrogen storage in metal hydrides involves chemically combining hydrogen gas with metal elements, forming metal hydrides. These offer significant safety benefits over gas and liquid storage methods, while also boasting higher hydrogen storage density (6.5 H atoms/cm3 for MgH2) as compared to hydrogen gas (0.99 H atoms/cm3) or liquid hydrogen (4.2 H atoms/cm3) 7. Hence, metal hydride storage is a safe, volume-efficient method suitable for onboard vehicle applications. Hydrogen typically stored within the lattice structure of metal hydrides is attached to metal atoms. Metal-hydrogen combinations yield two types of hydrides, $\alpha$-phase which absorbs some hydrogen, and $\beta$-phase where hydride is fully formed 8. Hydrogen storage in metal hydrides is influenced by various factors and involves several mechanistic steps. The ability of metals to dissociate hydrogen depends on surface structure, morphology, and purity 9. Factors such as the type of metal, type of hydride, and operating conditions impact the hydrogen storage capacity of metal hydrides. Light metals like Li, Be, Na, Mg, B, and Al form numerous metal–hydrogen compounds and are particularly intriguing owing to their lightweight and the large number of hydrogen atoms per metal atom, often in the order of H/M = 2 10, 11, 12, 13, 14.
@@
Many metals interact with hydrogen to form binary hydrides (MHn), but most binary hydrides do not present suitable hydrogen storage properties for serving as a hydrogen carrier. Some intermetallic compounds can also form hydrides with structural formulas of ABxHn, where element A with high hydrogen affinity is usually rare earth or alkaline earth metal forming a stable hydride, while element B has low hydrogen affinity and often forms only unstable hydrides 15, 4. The most extensively studied metal hydrides are AB2 and AB5 for hydrogen storage and fuel cell application 16. AB5-type intermetallic compounds are distinguished by their ease of activation, rapid kinetics of hydrogen absorption and desorption, and relatively high stability of hydrogen sorption properties during cyclic hydrogenation/dehydrogenation 17. Pressure-composition isotherms in H-AB5 systems show a single flat plateau with low H2 absorption-desorption hysteresis, which increases in substituted alloys 18. Ti and Zr are the most common A components in AB2-type compounds, and the B component is usually represented by a transition metal such as Mn, Cr, Fe, or V 17. AB2-type alloys are generally less easily activated than AB5 alloys, and they can be doped with small amounts ($\sim$1 at%) of rare earth elements to facilitate activation 19. They have excellent hydrogenation/dehydrogenation kinetics and cycle stability once activated 20. The AB2 components are less expensive than the AB5 components, but manufacturing of the AB2-type alloys presents some metallurgical challenges due to higher melting temperatures, high component reactivity, and other factors 20, 21. The most studied metal hydrides for hydrogen storage include sodium aluminum hydride, magnesium hydride, aluminum hydride, LaNi5, and ZrV2 22, 23, 18, 24.
@@
The introduction of data-driven materials design 25, 26, 27 has accelerated material discovery, processing, and manufacturing 28, 29, 30. Machine learning (ML) algorithms allow for the creation of new alloys based solely on previously collected data, either publicly available or reported in scientific literature. Metal hydrides are an ideal subject for machine learning algorithm research due to their diverse properties and broad range. Traditional methods for determining the optimized materials class and corresponding metal hydride composition based on desirable properties are difficult, time-consuming, and costly. ML learning techniques, on the other hand, enable rapid, productive, and efficient material class prediction for a specific hydrogen weight percent and operational conditions. Numerous groups have begun to use ML techniques to speed up this screening and gain more physical insight from massive amounts of data 31, 32, 33. Rezakazemi et al. used an adaptive neuro-fuzzy inference system to evaluate the performance of hydrogen-selective mixed matrix membranes under various operational conditions 34. Rahnama et al. predicted the hydrogen storage capacities in various metal hydrides using machine learning algorithms in another study and found that higher temperatures yielded higher hydrogen storage capacities 33, 35. Ahmed et al. recently predicted gravimetric and volumetric hydrogen capacities in metal-organic frameworks (MOFs) using ML algorithms 36. ML methods have been instrumental in predicting various properties of (MOFs) and even designing new structures. This approach has significantly changed the research landscape in this field, allowing for the rapid identification of promising MOF candidates 37, 38 for various applications, including hydrogen storage. In a recent wor, Sun et al. introduced a method that combines meta-learning and high-throughput molecular simulations to predict hydrogen storage in nanoporous materials efficiently. 39
@@
This study proposes a two-stage model for predicting hydrogen storage capacity in metal hydrides. The first stage involves unsupervised learning to extract latent features from the input data, and the second stage employs a multi-layer perceptron (MLP) trained to predict target properties. We developed a deep neural network-based model that can predict the hydrogen storage capacities in metal hydrides. We also identified new hydrogen storage materials using both feature-based approaches as well as using a large language model. The hydrogen storage capacity of these predicted materials was calculated using the above-mentioned approach and further compared with the results from the DFT-based methods. A chemical bonding analysis was performed to comprehend the relationship between storage capacity and the atomic environment, providing insights into the chemical bonds in the materials and their contribution to hydrogen storage capacity.
@@ 2 Methods
The architecture of our method is shown in Fig. 1, in a schematic way. We use a hybrid approach that has two parts: in the unsupervised learning part, an autoencoder is trained on the entire dataset, which consists of only unlabeled data, to learn a compact representation of the data in the bottleneck layer 40. A Multi-Layer Perceptron (MLP) is used in the supervised learning part to predict the hydrogen storage capacity. The major use of such an approach is to address the challenges of data scarcity and high feature dimensionality. In this study, the initial feature space has a high dimensionality and the available dataset is relatively small. As there are not too many machine learning studies available in the literature, it is not easy to choose proper descriptors. We have constructed 36 features using the Mendeleev software 41. To handle high-dimensional feature spaces with small datasets, we used an autoencoder to reduce the dimensionality of the data.
@@ 2.1 Unsupervised learning
Autoencoders are a kind of neural network that is used to extract features in an unsupervised manner. They are divided into two parts: an encoder, which compresses the input data into a lower-dimensional representation (also known as the latent representation or bottleneck), and a decoder, which reconstructs the original input from the compressed form 42.
@@
where $\sigma$ is the activation function, $W_{\phi}$ and $b_{\phi}$ are the weight and bias parameters of the encoder, respectively. For the activation function, we have used the RelU function as implemented in PyTorch 43.
@@
The goal of the autoencoder is to minimize the reconstruction error, defined as the difference between the original input $x$ and its reconstruction $\hat{x}$:
$L(\phi,\theta)=||\vec{x}-\hat{x}||^{2}$ (3)
This loss function was minimized using the Adam optimization method to find the optimal values for the parameters $\phi$ and $\theta$ say $\phi^{*}$ and $\theta^{*}$ 44. Once the autoencoder is trained, the encoder function $z=g_{\phi^{*}}(x)$ can be used to extract features from the input data by using the activations of the lower-dimensional representation $z$ as the features.
The network architecture for both encoder and decoder includes two linear layers and dropout layers for preventing overfitting. The input to the autoencoder is first passed through a linear layer with 256 units, followed by a Leaky ReLU activation function 45. This is then followed by a dropout layer with a dropout probability of 0.3 to provide regularization.
@@ 2.2 Supervised learning
Let $z_{i}$ be the input data, and let $y_{i}$ be the corresponding target (true) value. $N$ is the total number of data points and $L$ is the number of labeled data points.
Here we train a regressor, such as a Multilayer Perceptron (MLP), on the labeled data using the extracted features through the above-mentioned method as inputs. The MLP takes the latent features as input and predicts the target values based on the labeled data. The Loss function to minimize can be represented as:
$L=\frac{1}{N}\sum_{i=1}^{L}(y_{i}-\hat{y}_{i})^{2}$ (4)
Here $\hat{y}_{i}$ is the prediction of the MLP, and N is the total number of data points.
@@
First-principles calculations were conducted employing the density functional theory (DFT) within the Vienna Ab initio Software Package (VASP) 46, 47. To determine the electronic energy, projected augmented wave (PAW) potentials and plane-wave basis sets were utilized for both core and valence electrons, employing the Perdew–Burke–Ernzerhof (PBE) functional 48, 49. The calculations accounted for long-range interactions using the semi-empirical Grimme D2 dispersion method and incorporated non-spherical contributions to the PAW potentials within the code 50. A cutoff energy of 500 eV ensured convergence of all energies. Structural optimization employed the conjugate gradient algorithm 51, with convergence criteria set at 10-7 eV for energy and 0.01 eVÅ-1 for force. Spin polarization was enabled in all calculations, except for the isolated H2 closed shell molecule.
@@ 3 Dataset
For this study, we utilized the Hydrogen Storage Materials Database shared by the HyMARC Data Hub (https://datahub.hymarc.org/). This is an openly available database that can be accessed at https://datahub.hymarc.org/en/dataset/hydrogen-storage-materials-db. This database contains information on more than 2000 hydride materials and their properties. The Hydride Database classifies the hydrides into eight classes: A2B intermetallic compounds, AB intermetallic compounds, AB2 intermetallic compounds, AB5 intermetallic compounds, complex hydrides, Mg alloys, solid solution alloys, and Misc. In the dataset, more than 950 compounds have 1-2 wthydrogen storage capacity (Fig. 2). The average hydrogen storage capacity of all compounds is 2.1 wt%. The dataset has as high as 20.8 wthydrogen storage capacity complex hydrides, which is BeB4H8 52.
@@ 5.2 Case I: Using the MLP Directly for All the Features
For our primary experiment, the MLP was trained and tested using all of the 36 features derived from the Mendeleev software. The dataset was strategically partitioned into training and testing subsets, with respective ratios of 0.8 and 0.2. The resultant outcomes of this methodology are visually represented in Fig.4. The predictive efficacy, quantified by the $R^{2}$ coefficient, was deemed unsatisfactory 53. Such results emphasize the inherent difficulties connected with managing datasets with a high number of dimensions, especially when the amount of data is small.

### v5

# A Hybrid Machine Learning Framework for Predicting Hydrogen Storage Capacities in Metal Hydrides: Unsupervised Feature Learning with Deep Neural Networks

@@ Abstract
In this study, we present a sophisticated hybrid machine-learning framework that significantly improves the accuracy of predicting hydrogen storage capacities in metal hydrides. This is a critical challenge due to the scarcity of experimental data and the complexity of high-dimensional feature spaces. Our approach employs the power of unsupervised learning through the use of a state-of-the-art autoencoder. This autoencoder is trained on elemental descriptors obtained from Mendeleev software, enabling the extraction of a meaningful and lower dimensional latent space from the input data. This latent representation serves as the basis for our deep multi-layer perceptron (MLP) model, which consists of five layers and shows good precision in predicting hydrogen storage capacities. Furthermore, our results show very good agreement with the results obtained with density functional theory (DFT). In addition to addressing the limitations caused by limited and unevenly distributed data in the field of hydrogen storage materials, we also focus on discovering new materials that show promising opportunities for hydrogen storage. These materials were identified using both feature-based approaches and predictions generated by a large language model (LLM). A significant highlight of this work is the discovery of new hydrogen storage materials using a LLM, with a selected subset subsequently validated through density functional theory (DFT) calculations. Finally, our investigation into the effectiveness of transferring weights from the autoencoder to the MLP, in addition to the latent features, suggests that while this strategy slightly improves model performance as indicated by a slightly higher R2 value and lower RMSE, it emphasizes the intricate challenge of adapting pre-trained weights for specific supervised tasks.
@@
Hydrogen storage, deep learning, density functional theory, autoencoders, large language models
@@ 1 Introduction
Climate change, energy security and the potential depletion of resources due to growing global population and technological advances represent pressing challenges for humanity. 1, 2. To address these issues, promoting renewable energy is crucial as hydrogen is becoming an essential energy source for mobile and stationary applications. Hydrogen not only reduces environmental damage, but also reduces dependence on imported oil for countries without natural resources 3. There are various storage technologies for hydrogen, including compressed gas, cryogenic liquids and solid fuels as a chemical or physical combination with other materials such as metal hydrides, complex hydrides and carbon materials or produced in vehicles by on-site methanol reforming 4, 5. Each of these methods possesses unique attributes for hydrogen storage. Storage by absorption as chemical compounds or by adsorption on carbon materials carries safety benefits, necessitating a form of conversion or energy input for hydrogen release. Considerable effort has been invested into new hydrogen-storage systems, including metal, chemical, or complex hydrides, and carbon nanostructures 6, 7.
@@
Hydrogen storage in metal hydrides involves chemically combining hydrogen gas with metal elements, forming metal hydrides. These offer significant safety benefits over gas and liquid storage methods, while also boasting higher hydrogen storage density (6.5 H atoms/cm3 for MgH2) as compared to hydrogen gas (0.99 H atoms/cm3) or liquid hydrogen (4.2 H atoms/cm3) 8. Hence, metal hydride storage is a safe, volume-efficient method suitable for onboard vehicle applications. Hydrogen typically stored within the lattice structure of metal hydrides is attached to metal atoms. Metal-hydrogen combinations yield two types of hydrides, $\alpha$-phase which absorbs some hydrogen, and $\beta$-phase where hydride is fully formed 9. Hydrogen storage in metal hydrides is influenced by various factors and involves several mechanistic steps. The ability of metals to dissociate hydrogen depends on surface structure, morphology, and purity 10. Factors such as the type of metal, type of hydride, and operating conditions impact the hydrogen storage capacity of metal hydrides. Light metals like Li, Be, Na, Mg, B, and Al form numerous metal–hydrogen compounds and are particularly intriguing owing to their lightweight and the large number of hydrogen atoms per metal atom, often in the order of H/M = 2 11, 12, 13, 14, 15.
@@
Many metals have the capacity to engage with hydrogen in order to produce binary hydrides (MHn). Nonetheless, the majority of binary hydrides lack the necessary properties for effective hydrogen storage as a carrier. Certain intermetallic compounds have the ability to generate hydrides with structural formulas of ABxHn. In this scenario, element A, which typically belongs to the rare earth or alkaline earth metal category due to its strong hydrogen affinity, forms a stable hydride, while element B tends to create only unstable hydrides given its low hydrogen affinity 16, 4. The metal hydrides that have been extensively examined are those falling under AB2 and AB5, particularly in relation to hydrogen storage and applications in fuel cells 17. Intermetallic compounds of the AB5 type stand out due to their easily achievable activation, swift kinetics in hydrogen absorption and release, as well as the relatively high stability exhibited in hydrogen sorption properties throughout cyclic hydrogenation/dehydrogenation processes 18. Pressure-composition isotherms in H-AB5 systems show a single flat plateau with low H2 absorption-desorption hysteresis, which increases in substituted alloys 19. Ti and Zr are the most common A components in AB2-type compounds, and the B component is usually represented by a transition metal such as Mn, Cr, Fe, or V 18. AB2-type alloys are generally less easily activated than AB5 alloys, and they can be doped with small amounts ($\sim$1 at%) of rare earth elements to facilitate activation 20. They have excellent hydrogenation/dehydrogenation kinetics and cycle stability once activated 21. The AB2 components are less expensive than the AB5 components, but manufacturing of the AB2-type alloys presents some metallurgical challenges due to higher melting temperatures, high component reactivity, and other factors 21, 22. The most studied metal hydrides for hydrogen storage include sodium aluminum hydride, magnesium hydride, aluminum hydride, LaNi5, and ZrV2 23, 24, 19, 25.
@@
The introduction of data-driven materials design 26, 27, 28 has accelerated material discovery, processing, and manufacturing 29, 30, 31, 32, 33. Machine learning (ML) algorithms allow for the creation of new alloys based solely on previously collected data, either publicly available or reported in scientific literature. Metal hydrides are an ideal subject for machine learning algorithm research due to their diverse properties and broad range. Traditional methods for determining the optimized materials class and corresponding metal hydride composition based on desirable properties are difficult, time-consuming, and costly. ML techniques, on the other hand, enable rapid, productive, and efficient material class prediction for a specific hydrogen weight percent and operational conditions. Numerous groups have begun to use ML techniques to speed up this screening and gain more physical insight from massive amounts of data 34, 35, 36. Rezakazemi et al. used an adaptive neuro-fuzzy inference system to evaluate the performance of hydrogen-selective mixed matrix membranes under various operational conditions 37. Rahnama et al. predicted the hydrogen storage capacities in various metal hydrides using machine learning algorithms in another study and found that higher temperatures yielded higher hydrogen storage capacities 36, 38. Ahmed et al. recently predicted gravimetric and volumetric hydrogen capacities in metal-organic frameworks (MOFs) using ML algorithms 39. ML methods have been instrumental in predicting various properties of (MOFs) and even designing new structures. This approach has significantly changed the research landscape in this field, allowing for the rapid identification of promising MOF candidates 40, 41 for various applications, including hydrogen storage. In a recent work, Sun et al. introduced a method that combines meta-learning and high-throughput molecular simulations to predict hydrogen storage in nanoporous materials efficiently. 42
@@
In this study, we propose a two-stage model for predicting hydrogen storage capacity in metal hydrides. The first stage involves unsupervised learning to extract latent features from the input data, and the second stage employs a multi-layer perceptron (MLP) trained to predict target properties. We developed a deep neural network-based model that can predict the hydrogen storage capacities in metal hydrides. We also identified new hydrogen storage materials using both feature-based approaches as well as using a large language model (LLM). The hydrogen storage capacity of these predicted materials was calculated using the above-mentioned approach and further compared with the results from the DFT-based methods. A chemical bonding analysis was performed to comprehend the relationship between storage capacity and the atomic environment, providing insights into the chemical bonds in the materials and their contribution to hydrogen storage capacity.
It is worth noting that the application of LLMs in materials science has shown significant potential as they leverage their advanced natural language processing capabilities to predict and design novel materials. Recent advances include the development of the Materials Informatics Transformer (MatInFormer), which leverages tokenization of crystallographic space group information for high-precision predictions of material properties, particularly in metal-organic frameworks (MOFs). 43. Furthermore, transformer models such as GPT and BART have been successfully applied to generative design tasks, producing chemically valid and novel material compositions with a high degree of accuracy in terms of charge neutrality and electronegativity balance 44. As another example, the HoneyBee model illustrates the specific adaptation of LLMs for materials science through iterative fine-tuning processes 45 and demonstrates superior performance on domain-specific tasks.
@@ 2 Methods
The architecture of our method is shown in Fig. 1, in a schematic way. We use a hybrid approach that has two parts: in the unsupervised learning part, an autoencoder is trained on the entire dataset, which consists of only unlabeled data, to learn a compact representation of the data in the bottleneck layer 46. A Multi-Layer Perceptron (MLP) is used in the supervised learning part to predict the hydrogen storage capacity. The major use of such an approach is to address the challenges of data scarcity and high feature dimensionality. In this study, the initial feature space has a high dimensionality and the available dataset is relatively small. As there are not too many machine learning studies available in the literature, it is not easy to choose proper descriptors. We have constructed 36 features using the Mendeleev software 47. To handle high-dimensional feature spaces with small datasets, we used an autoencoder to reduce the dimensionality of the data.
@@ 2.1 Unsupervised learning
Autoencoders are a kind of neural network that is used to extract features in an unsupervised manner. They are divided into two parts: an encoder, which compresses the input data into a lower-dimensional representation (also known as the latent representation or bottleneck), and a decoder, which reconstructs the original input from the compressed form 48.
@@
where $\sigma$ is the activation function, $W_{\phi}$ and $b_{\phi}$ are the weight and bias parameters of the encoder, respectively. For the activation function, we have used the RelU function as implemented in PyTorch 49.
@@
The goal of the autoencoder is to minimize the reconstruction error, defined as the difference between the original input $x$ and its reconstruction $\hat{x}$,
$\mathcal{L}(\phi,\theta)=||\vec{x}-\hat{x}||^{2}$ (3)
This loss function was minimized using the Adam optimization method to find the optimal values for the parameters $\phi$ and $\theta$ say $\phi^{*}$ and $\theta^{*}$ 50. Once the autoencoder is trained, the encoder function $z=g_{\phi^{*}}(x)$ can be used to extract features from the input data by using the activations of the lower-dimensional representation $z$ as the features.
The network architecture for both the encoder and decoder includes two linear layers and dropout layers to prevent overfitting. The input to the autoencoder is first passed through a linear layer with 256 units, followed by a Leaky ReLU activation function 51. This is then followed by a dropout layer with a dropout probability of 0.3 to provide regularization.
@@ 2.2 Supervised learning
Let $z_{i}$ be the input data, and let $y_{i}$ be the corresponding target (true) value, and $L$ is the number of labeled data points.
We train a regressor, such as a Multilayer Perceptron (MLP), on the labeled data using the extracted features through the aforementioned method as inputs. The MLP takes these latent features as input and predicts the target values based on the labeled data. The loss function to minimize can be expressed as:
$\mathcal{L}=\frac{1}{L}\sum_{i=1}^{L}(y_{i}-\hat{y}_{i})^{2}$ (4)
Here, $\hat{y}_{i}$ represents the predicted values from the MLP, and $\mathcal{L}$ is the mean squared error loss over the labeled data points.
@@
First-principles calculations were conducted employing the density functional theory (DFT) within the Vienna Ab initio Software Package (VASP) 52, 53. To determine the electronic energy, projected augmented wave (PAW) potentials and plane-wave basis sets were utilized for both core and valence electrons, employing the Perdew–Burke–Ernzerhof (PBE) functional 54, 55. The calculations accounted for long-range interactions using the semi-empirical Grimme D2 dispersion method and incorporated non-spherical contributions to the PAW potentials within the code 56. A cutoff energy of 500 eV ensured convergence of all energies. Structural optimization employed the conjugate gradient algorithm 57, with convergence criteria set at 10-7 eV for energy and 0.01 eVÅ-1 for force. Spin polarization was enabled in all calculations, except for the isolated H2 closed shell molecule.
@@ 3 Dataset
For this study, we utilized the Hydrogen Storage Materials Database shared by the HyMARC Data Hub (https://datahub.hymarc.org/). This is an openly available database that can be accessed at https://datahub.hymarc.org/en/dataset/hydrogen-storage-materials-db. This database contains information on more than 2000 hydride materials and their properties. The Hydride Database classifies the hydrides into eight classes: A2B intermetallic compounds, AB intermetallic compounds, AB2 intermetallic compounds, AB5 intermetallic compounds, complex hydrides, Mg alloys, solid solution alloys, and Misc. In the dataset, more than 950 compounds have 1-2 wthydrogen storage capacity (Fig. 2). The average hydrogen storage capacity of all compounds is 2.1 wt%. The dataset has as high as 20.8 wthydrogen storage capacity complex hydrides, which is BeB4H8 58.
@@ 5.2 Case I: Using the MLP Directly for All the Features
For our primary experiment, the MLP was trained and tested using all of the 36 features derived from the Mendeleev software. The dataset was strategically partitioned into training and testing subsets, with respective ratios of 0.8 and 0.2. The resultant outcomes of this methodology are visually represented in Fig.4. The predictive efficacy, quantified by the $R^{2}$ coefficient, was deemed unsatisfactory. The coefficient of determination ($R^{2}$) is 0.764, though a reasonably good fit, indicating that only approximately 76.4% of the variance in the true values can be explained by the model. Such results emphasize the inherent difficulties connected with managing datasets with a high number of dimensions, especially when the amount of data is small.